Abbey and Mia are in the basement playing pool. On Abbey's recent shot, the cue ball was moving east at 82 cm/s when it struck the slower 5-ball moving in the same direction at 24 cm/s. The 5-ball immediately speeds up to 52 cm/s. Determine the post-collision speed of the cue ball in cm/s.


1
Expert's answer
2022-04-17T17:07:24-0400

"m_1=m_2,"

"v_1+v_2=u_1+u_2,"

"u_1=v_1+v_2-u_2=54~\frac{cm}s ."